# Diffusion Models in Simulation-Based Inference: A Tutorial Review

### v1

@@ 4.1.3 Inference-Time Prior Adaptation
In many applications, priors can change after the model has been trained. For instance, one can shift from population-level to subject-specific priors, or perform Bayesian updating, which promotes yesterday’s posterior to today’s prior. Building on the guidance view above, prior adaptation handles the case where only the prior changes while the observation model remains the same. Instead of retraining, the learned posterior score approximator can be reused by adding an inference-time correction that propagates the prior ratio through diffusion time [yang2025priorguide].
@@
The first term is the original posterior score under the training prior, while the second term serves as a guidance correction that propagates the effect of the new prior at diffusion time $t$. To evaluate the additional term in practice, [yang2025priorguide] propose to approximate the reverse transition kernel $p(\boldsymbol{\theta}_{0}\mid\boldsymbol{\theta}_{t},\mathbf{y}_{\text{obs}})$ and the prior ratio using a generalized mixture of Gaussians, which yields a tractable closed form correction.
Furthermore, prior adaptation could unlock continual Bayesian updating by using the posterior from one step as the new prior in the next, thereby chaining inference tasks without retraining the diffusion estimator. It also connects to validation methods for neural posterior estimators. For instance, simulation-based calibration (SBC) [cook2006validation, talts2018validating, sailynoja2022graphical] provides a default framework to assess calibration properties (e.g., over- or underdispersion) of the estimator in silico, and posterior SBC extends this idea by conditioning on a fixed observation $\mathbf{y}_{\text{obs}}$ [sailynoja2025posterior]. This would usually require training a model to handle an augmented posterior $p(\boldsymbol{\theta}\mid\mathbf{y},\mathbf{y}_{\text{obs}})$. Under the prior adaptation view, however, this is either equivalent to changing the prior to $p(\boldsymbol{\theta}\mid\mathbf{y}_{\text{obs}})$ or to composing two observations, $\mathbf{y}$ and $\mathbf{y}_{\text{obs}}$. Both variants can be implemented with diffusion models without retraining on both observations simultaneously. allowing posterior SBC to be efficiently checked.
@@ 4.2 Aggregating Scores for Structured Observations
Such structures arise widely across scientific domains. Examples include: repeated measurements in cognitive psychology [lee2011cognitive]; single-cell or multi-modal biological data with large i.i.d. cohorts [kilian2024longitudinal]; dynamical systems and ODE/SDE models in epidemiology or biology [radev2021outbreakflow]; multi-level or grouped observations in pharmacokinetics [wakefield1996pharmacokinetic]; or sequential updating in climate modeling [houtekamer2016review], to name just a few.
Score decomposition converts these probabilistic symmetries into additive components that can be estimated independently and composed at inference time, allowing training on partial simulations rather than full-model simulation. As a result, the choice of a Bayesian model determines whether parameters are inferred for each observation independently or shared across many observations. Thus, common pooling assumptions in Bayesian analysis [bda3] shape how diffusion models can exploit factorizations to reduce simulation and training costs dramatically.
@@ 4.2.1 No-Pooling Models
Consequently, no information is shared between observations, and each posterior estimate relies solely on its corresponding individual observation. Most diffusion model adaptations in SBI focus on the no-pooling setting (LABEL:tab:papers), which has emerged as the predominant target of SBI applications [zammit2024neural] and benchmarks [lueckmann2021benchmarking]. When the number of observations $R$ is large, a single diffusion model can be trained to estimate all posteriors for $\{\mathbf{y}^{(r)}_{\text{obs}}\}^{R}_{r=1}$ without further re-training (i.e., amortized inference, see Section 2.1).
@@
So far, we have not paid much attention to the structure of each $\mathbf{y}$. Many Bayesian models assume exchangeable observations $\mathbf{y}_{1:N}:=\{\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{y}_{N}\}\equiv\mathbf{y}_{\text{obs}}$, for instance, due to repeated independent experiments, leading to a factorizable model [gelman1995bayesian]:
@@
In NPE, one typically uses a permutation-invariant network, such as DeepSet [zaheer2017deep] or SetTransformer [lee2019set], to encode this exchangeability [radev2020bayesflow]. However, this approach requires $N$ simulator calls to generate a single training instance. With score-based models, simulation and training efficiency can be increased by targeting the compositional posterior
@@
which follows directly from Bayes’ rule. Transforming this expression to score-space, [geffner2023compositional] introduced compositional score matching (CSM) and showed that the time-dependent score of the total posterior can be expressed as a linear combination of the prior and $N$ posterior factors:
@@
However, naive compositional estimation is prone to error accumulation, making (29) unstable and necessitating careful scheduling or error damping [arruda2025compositional]. Moreover, instead of composing over all $N$ elements, one can also use partial factorization over multiple elements [geffner2023compositional], which can reduce error accumulation but increases the computational costs during training as more simulations are needed.
@@
Ordered observations (i.e., time series) $\mathbf{y}_{1:N}:=(\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{y}_{N})\equiv\mathbf{y}_{\text{obs}}$ are ubiquitous in Bayesian modeling. Practically relevant examples include ODE/SDE type models such as susceptible-infected-recovered (SIR) models [radev2021outbreakflow], or Kolmogorov flows [Rozet2023-dataassi]. A common case is a first-order Markov model, where dependencies are limited to adjacent observations in a sequence:
@@
Compositional estimation can be extended to Markov models by learning a local score estimator for $\nabla_{\boldsymbol{\theta}_{t}}\!\log p(\boldsymbol{\theta}_{t}\mid\mathbf{y}_{n},\mathbf{y}_{n-1})$ which requires only single-step transitions from $\mathbf{y}_{n-1}$ to $\mathbf{y}_{n}$ during training [gloeckler2025compositional]. This avoids the need to evolve a first-order Markov process over $N$ steps for each training instance. At inference, the individual scores can be aggregated similarly to (29) across the time series:
$\nabla_{\boldsymbol{\theta}_{t}}\!\log p(\boldsymbol{\theta}_{t}\mid\mathbf{y}_{1:N})\approx\,$ $\Lambda(\boldsymbol{\theta}_{t})^{-1}\Bigl((1-N)\Sigma^{-1}_{t}\,\nabla_{\boldsymbol{\theta}_{t}}\!\log p(\boldsymbol{\theta}_{t})$
$\,+\Sigma^{-1}_{t,1}\,\nabla_{\boldsymbol{\theta}_{t}}\!\log p(\boldsymbol{\theta}_{t}\mid\mathbf{y}_{1})+\sum_{n=2}^{N}\Sigma^{-1}_{t,n}\,\nabla_{\boldsymbol{\theta}_{t}}\!\log p(\boldsymbol{\theta}_{t}\mid\mathbf{y}_{n},\mathbf{y}_{n-1})\Bigr),$ (31)
where $\Lambda(\boldsymbol{\theta}_{t})=\sum_{n=1}^{N}\Sigma^{-1}_{t,n}+(1-N)\Sigma_{t}^{-1}$, and $\Sigma_{t}^{-1}$ and $\Sigma^{-1}_{t,n}$ are the denoising prior and posterior precision matrices respectively. The marginal prior score $\nabla_{\boldsymbol{\theta}_{t}}\!\log p(\boldsymbol{\theta}_{t})$ is typically known analytically, while [linhart2024diffusion] propose two approximation methods for $\Sigma^{-1}_{t,n}$: GAUSS, which assumes that the posterior is Gaussian, and analytically computes the denoising posterior covariances; and JAC, which iteratively estimates them via Tweedie’s moment projection [boys2024tweedie] using the Jacobian of the score model. Ensuring $\Lambda(\boldsymbol{\theta}_{t})$ to be positive definite makes the approximation more stable [gloeckler2025compositional], but [touron2025erroranalysiscompositionalscorebased] show that error accumulation under the assumptions of GAUSS is directly related to the sampling quality of each individual posterior. A key ingredient here is a proposal distribution $\mathbf{y}_{n}\sim q_{n}(\mathbf{y})$ for generating single-step transitions at arbitrary time points $n$ without simulating the full-time horizon. Crucially, the support of the proposal should overlap with the support of the likelihood. If the proposal is too narrow, important regions of the transition space may never be explored; if it is too wide, the estimator may waste capacity on implausible transitions and learn slowly.
@@
In recursive Bayesian estimation, the aim is to estimate the current latent state $\boldsymbol{\theta}_{n}$ given a sequence of past observations $\mathbf{y}_{1:n}$. This is a standard setting in filtering problems, where sequential Bayesian updates are carried out in high-dimensional dynamical systems using ensemble-based approximations such as the Ensemble Kalman Filter [evensen2009data, houtekamer2016review]. The posterior distribution of interest is therefore
@@
For every new observation, the posterior needs to be updated, hence both the prior $p(\boldsymbol{\theta}_{n}\mid\mathbf{y}_{1:n-1})$ and the posterior are time-varying. Similar to (18), applying the score operator to Bayes’ rule decomposes this modeling problem into the following [Andry2023-dataAssimilationSimulation-basedInferences]:
@@
This shows that the posterior score is the sum of the prior score, which describes the dynamics of the latent state, and the perturbed likelihood score, which incorporates the information from the observation process. One can either train a score network directly to approximate the posterior score on the left (Section 3), or learn the two terms on the right separately and combine them at inference time (Section 4.1). The second option can be advantageous if one has access to an explicit description of the observation process, as this allows the likelihood score to be approximated directly. This line of work is more closely related to inverse problems rather than SBI, and has been applied to posterior sampling in noisy inverse problems [Chung2022-diffusionPosteriorSamplingGeneralNoisyInverseProblemse]. In this way, decomposing the posterior clarifies that posterior inference in state-space models with diffusion can be addressed either by directly learning the posterior score or by combining learned dynamics with known likelihood terms [Andry2023-dataAssimilationSimulation-basedInferences].
@@ 4.2.2 Complete Pooling Models
that aggregates all observations without considering potential group or individual differences, thus enforcing complete homogeneity across data points [Bardenet2017talldata]. Complete pooling is a rare target in SBI (and Bayesian inference in general) since its homogeneity assumption is ordinarily too restrictive in practice. However, it provides a straightforward probabilistic aggregation recipe that does not rely on ad hoc postprocessing. As shown by [linhart2024diffusion], diffusion models can pool data from multiple experiments by using compositional score aggregation.
The internal probabilistic symmetry of each $\mathbf{y}_{\text{obs}}^{(r)}$ is problem-dependent (e.g., exchangeable, Markovian, etc.) and thus, each $\mathbf{y}_{\text{obs}}^{(r)}$ can be a sequence or a set of varying length $N_{r}$. Training a standard diffusion model to sample directly from the posterior over all observations (34) requires $\sum N_{r}$ simulator runs per training instance if $\mathbf{y}_{\text{obs}}^{(r)}$ itself consists of multiple elements, which is computationally prohibitive. Thus, [linhart2024diffusion] proposed to estimate the individual posterior scores for each $\mathbf{y}_{\text{obs}}^{(r)}$ and combine them using the approximate compositional score expression
$\nabla_{\boldsymbol{\theta}_{t}}\!\log p(\boldsymbol{\theta}_{t}\mid\{\mathbf{y}_{\text{obs}}^{(r)}\}_{r=1}^{R})\approx\Lambda(\boldsymbol{\theta}_{t})^{-1}\left((1-R)\Sigma^{-1}_{t}\,\nabla_{\boldsymbol{\theta}_{t}}\!\log p(\boldsymbol{\theta}_{t})+\sum_{r=1}^{R}\Sigma^{-1}_{t,r}\,\nabla_{\boldsymbol{\theta}_{t}}\!\log p(\boldsymbol{\theta}_{t}\mid\mathbf{y}_{\text{obs}}^{(r)})\right),$ (35)
@@ 4.2.3 Partial Pooling Models
A partial pooling or hierarchical model represents an intermediate between complete pooling and no-pooling [gelman2012we]. In these models, not only do we have structured observations, but we also introduce additional structure to the parameters: observation-specific parameters $\boldsymbol{\theta}^{(r)}$, which are assumed to be drawn from a shared higher-level distribution governed by hyperparameters $\boldsymbol{\eta}$. This structure naturally arises in many domains with inherent hierarchies, such as patients in pharmacokinetic models [wakefield1996pharmacokinetic], longitudinal patient omics [kilian2024longitudinal], single-cells in gene expression models [llamosi2016population], or participants in trails in cognitive modeling [lee2011cognitive].
@@
Such models allow individual parameters to be informed by global estimates (38), stabilizing inference at low $N_{r}$ by borrowing information across observations. Even though partial pooling has been espoused as a default choice in Bayesian analysis [gelman2012we, mcelreath2018statistical], it typically incurs high computational cost and has only recently received attention in SBI [heinrichh2024hierarchical, habermann2024amortized, arruda2024amortized, charles2025tokenised].
To amortize two-level models, for example, we can train two diffusion models, one for the global and one for the local posterior. During training, the local model is conditioned on the true parameter $\boldsymbol{\eta}$ and during inference, an ancestral sampling scheme can be used by sampling first from the global and then from the local model. Both models can share an embedding for individual observations. However, targeting the joint posterior (36) or even the global posterior (37) without composition would require $R$ simulations per training instance. Thus, [arruda2025compositional] ported compositional score estimation to the hierarchical setting, enabling amortized hierarchical estimation without ever simulating the full hierarchical model. To address stability issues, the authors applied an error-damping schedule $d(t)$ to the reverse diffusion with a mini-batch sampler using $M\ll S$ randomly sampled observations at each time step $t$. The resulting compositional estimator,
$\nabla_{\boldsymbol{\eta}_{t}}\!\log p(\boldsymbol{\eta}_{t}\mid\{\mathbf{y}_{\text{obs}}^{(r)}\}_{r=1}^{R})=d(t)\cdot\bigg((1-R)(1-t)\nabla_{\boldsymbol{\eta}}\!\log p(\boldsymbol{\eta})\bigl\rvert_{\boldsymbol{\eta}=\boldsymbol{\eta}_{t}}+\frac{R}{M}\sum_{m=1}^{M}\nabla_{\boldsymbol{\eta}_{t}}\!\log p(\boldsymbol{\eta}_{t}\mid\mathbf{y}_{\text{obs}}^{(m)})\bigg),$ (39)

### v2

@@ 4.1.3 Inference-Time Prior Adaptation
In many applications, priors can change after the model has been trained. For instance, one can shift from population-level to subject-specific priors, or perform Bayesian updating, which promotes yesterday’s posterior to today’s prior. Building on the guidance view above, prior adaptation handles the case where only the prior changes while the observation model remains the same. Instead of retraining, the learned posterior score approximator can be reused by adding an inference-time correction that propagates the prior ratio through diffusion time [192].
@@
The first term is the original posterior score under the training prior, while the second term serves as a guidance correction that propagates the effect of the new prior at diffusion time $t$. To evaluate the additional term in practice, [192] propose to approximate the reverse transition kernel $p(\boldsymbol{\theta}_{0}\mid\boldsymbol{\theta}_{t},\mathbf{y}_{\text{obs}})$ and the prior ratio using a generalized mixture of Gaussians, which yields a tractable closed-form correction.
Furthermore, prior adaptation could unlock continual Bayesian updating by using the posterior from one step as the new prior in the next, thereby chaining inference tasks without retraining the diffusion estimator. It also connects to validation methods for neural posterior estimators. For instance, simulation-based calibration (SBC) [30, 169, 145] provides a default framework to assess calibration properties (e.g., over- or underdispersion) of the estimator in silico, and posterior SBC extends this idea by conditioning on a fixed observation $\mathbf{y}_{\text{obs}}$ [146]. This would usually require training a model to handle an augmented posterior $p(\boldsymbol{\theta}\mid\mathbf{y},\mathbf{y}_{\text{obs}})$. Under the prior adaptation view, however, this is either equivalent to changing the prior to $p(\boldsymbol{\theta}\mid\mathbf{y}_{\text{obs}})$ or to composing two observations, $\mathbf{y}$ and $\mathbf{y}_{\text{obs}}$. Both variants can be implemented with diffusion models without retraining on both observations simultaneously. allowing posterior SBC to be efficiently checked.
@@ 4.2 Aggregating Scores for Structured Observations
Such structures arise widely across scientific domains. Examples include: repeated measurements in cognitive psychology [100]; single-cell or multi-modal biological data with large i.i.d. cohorts [91]; dynamical systems and ODE/SDE models in epidemiology or biology [138]; multi-level or grouped observations in pharmacokinetics [185]; or sequential updating in climate modeling [79], to name just a few.
Score decomposition converts these probabilistic symmetries into additive components that can be estimated independently and composed at inference time, allowing training on partial simulations rather than full-model simulation. As a result, the choice of a Bayesian model determines whether parameters are inferred for each observation independently or shared across many observations. Thus, common pooling assumptions in Bayesian analysis [54] shape how diffusion models can exploit factorizations to reduce simulation and training costs dramatically.
@@ 4.2.1 No-Pooling Models
Consequently, no information is shared between observations, and each posterior estimate relies solely on its corresponding individual observation. Most diffusion model adaptations in SBI focus on the no-pooling setting (LABEL:tab:papers), which has emerged as the predominant target of SBI applications [197] and benchmarks [114]. When the number of observations $R$ is large, a single diffusion model can be trained to estimate all posteriors for $\{\mathbf{y}^{(r)}_{\text{obs}}\}^{R}_{r=1}$ without further re-training (i.e., amortized inference, see Section 2.1).
@@
So far, we have not paid much attention to the structure of each $\mathbf{y}$. Many Bayesian models assume exchangeable observations $\mathbf{y}_{1:N}:=\{\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{y}_{N}\}\equiv\mathbf{y}_{\text{obs}}$, for instance, due to repeated independent experiments, leading to a factorizable model [55]:
@@
In NPE, one typically uses a permutation-invariant network, such as DeepSet [196] or SetTransformer [99], to encode this exchangeability [139]. However, this approach requires $N$ simulator calls to generate a single training instance. With score-based models, simulation and training efficiency can be increased by targeting the compositional posterior
@@
which follows directly from Bayes’ rule. Transforming this expression to score-space, [53] introduced compositional score matching (CSM) and showed that the time-dependent score of the total posterior can be expressed as a linear combination of the prior and $N$ posterior factors:
@@
However, naive compositional estimation is prone to error accumulation, making (29) unstable and necessitating careful scheduling or error damping [4]. Moreover, instead of composing over all $N$ elements, one can also use partial factorization over multiple elements [53], which can reduce error accumulation but increases the computational costs during training as more simulations are needed.
@@
Ordered observations (i.e., time series) $\mathbf{y}_{1:N}:=(\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{y}_{N})\equiv\mathbf{y}_{\text{obs}}$ are ubiquitous in Bayesian modeling. Practically relevant examples include ODE/SDE type models such as susceptible-infected-recovered (SIR) models [138], or Kolmogorov flows [143]. A common case is a first-order Markov model, where dependencies are limited to adjacent observations in a sequence:
@@
Compositional estimation can be extended to Markov models by learning a local score estimator for $\nabla_{\boldsymbol{\theta}_{t}}\!\log p(\boldsymbol{\theta}_{t}\mid\mathbf{y}_{n},\mathbf{y}_{n-1})$ which requires only single-step transitions from $\mathbf{y}_{n-1}$ to $\mathbf{y}_{n}$ during training [59]. This avoids the need to evolve a first-order Markov process over $N$ steps for each training instance. At inference, the individual scores can be aggregated similarly to (29) across the time series:
$\begin{split}\nabla_{\boldsymbol{\theta}_{t}}\!\log p(\boldsymbol{\theta}_{t}\mid\mathbf{y}_{1:N})\approx&\,\Lambda(\boldsymbol{\theta}_{t})^{-1}\Bigl((1-N)\Sigma^{-1}_{t}\,\nabla_{\boldsymbol{\theta}_{t}}\!\log p(\boldsymbol{\theta}_{t})\\ &+\Sigma^{-1}_{t,1}\,\nabla_{\boldsymbol{\theta}_{t}}\!\log p(\boldsymbol{\theta}_{t}\mid\mathbf{y}_{1})+\sum_{n=2}^{N}\Sigma^{-1}_{t,n}\,\nabla_{\boldsymbol{\theta}_{t}}\!\log p(\boldsymbol{\theta}_{t}\mid\mathbf{y}_{n},\mathbf{y}_{n-1})\Bigr),\end{split}$
where $\Lambda(\boldsymbol{\theta}_{t})=\sum_{n=1}^{N}\Sigma^{-1}_{t,n}+(1-N)\Sigma_{t}^{-1}$, and $\Sigma_{t}^{-1}$ and $\Sigma^{-1}_{t,n}$ are the denoising prior and posterior precision matrices respectively. The marginal prior score $\nabla_{\boldsymbol{\theta}_{t}}\!\log p(\boldsymbol{\theta}_{t})$ is typically known analytically, while [105] propose two approximation methods for $\Sigma^{-1}_{t,n}$: GAUSS, which assumes that the posterior is Gaussian, and analytically computes the denoising posterior covariances; and JAC, which iteratively estimates them via Tweedie’s moment projection [17] using the Jacobian of the score model. Ensuring $\Lambda(\boldsymbol{\theta}_{t})$ to be positive definite makes the approximation more stable [59], but [176] show that error accumulation under the assumptions of GAUSS is directly related to the sampling quality of each individual posterior. A key ingredient here is a proposal distribution $\mathbf{y}_{n}\sim q_{n}(\mathbf{y})$ for generating single-step transitions at arbitrary time points $n$ without simulating the full-time horizon. Crucially, the support of the proposal should overlap with the support of the likelihood. If the proposal is too narrow, important regions of the transition space may never be explored; if it is too wide, the estimator may waste capacity on implausible transitions and learn slowly.
@@
In recursive Bayesian estimation, the aim is to estimate the current latent state $\boldsymbol{\theta}_{n}$ given a sequence of past observations $\mathbf{y}_{1:n}$. This is a standard setting in filtering problems, where sequential Bayesian updates are carried out in high-dimensional dynamical systems using ensemble-based approximations such as the Ensemble Kalman Filter [44, 79]. The posterior distribution of interest is therefore
@@
For every new observation, the posterior needs to be updated, hence both the prior $p(\boldsymbol{\theta}_{n}\mid\mathbf{y}_{1:n-1})$ and the posterior are time-varying. Similar to (18), applying the score operator to Bayes’ rule decomposes this modeling problem into the following [3]:
@@
This shows that the posterior score is the sum of the prior score, which describes the dynamics of the latent state, and the perturbed likelihood score, which incorporates the information from the observation process. One can either train a score network directly to approximate the posterior score on the left (Section 3), or learn the two terms on the right separately and combine them at inference time (Section 4.1). The second option can be advantageous if one has access to an explicit description of the observation process, as this allows the likelihood score to be approximated directly. This line of work is more closely related to inverse problems rather than SBI, and has been applied to posterior sampling in noisy inverse problems [28]. In this way, decomposing the posterior clarifies that posterior inference in state-space models with diffusion can be addressed either by directly learning the posterior score or by combining learned dynamics with known likelihood terms [3].
@@ 4.2.2 Complete Pooling Models
that aggregates all observations without considering potential group or individual differences, thus enforcing complete homogeneity across data points [12]. Complete pooling is a rare target in SBI (and Bayesian inference in general) since its homogeneity assumption is ordinarily too restrictive in practice. However, it provides a straightforward probabilistic aggregation recipe that does not rely on ad hoc postprocessing. As shown by [105], diffusion models can pool data from multiple experiments by using compositional score aggregation.
The internal probabilistic symmetry of each $\mathbf{y}_{\text{obs}}^{(r)}$ is problem-dependent (e.g., exchangeable, Markovian, etc.) and thus, each $\mathbf{y}_{\text{obs}}^{(r)}$ can be a sequence or a set of varying length $N_{r}$. Training a standard diffusion model to sample directly from the posterior over all observations (33) requires $\sum N_{r}$ simulator runs per training instance if $\mathbf{y}_{\text{obs}}^{(r)}$ itself consists of multiple elements, which is computationally prohibitive. Thus, [105] proposed to estimate the individual posterior scores for each $\mathbf{y}_{\text{obs}}^{(r)}$ and combine them using the approximate compositional score expression
$\begin{split}\nabla_{\boldsymbol{\theta}_{t}}\!\log p(\boldsymbol{\theta}_{t}\mid\{\mathbf{y}_{\text{obs}}^{(r)}\}_{r=1}^{R})\approx&\,\Lambda(\boldsymbol{\theta}_{t})^{-1}\Bigl((1-R)\Sigma^{-1}_{t}\,\nabla_{\boldsymbol{\theta}_{t}}\!\log p(\boldsymbol{\theta}_{t})\\ &+\sum_{r=1}^{R}\Sigma^{-1}_{t,r}\,\nabla_{\boldsymbol{\theta}_{t}}\!\log p(\boldsymbol{\theta}_{t}\mid\mathbf{y}_{\text{obs}}^{(r)})\Bigr),\end{split}$ (34)
@@ 4.2.3 Partial Pooling Models
A partial pooling or hierarchical model represents an intermediate between complete pooling and no-pooling [56]. In these models, not only do we have structured observations, but we also introduce additional structure to the parameters: observation-specific parameters $\boldsymbol{\theta}^{(r)}$, which are assumed to be drawn from a shared higher-level distribution governed by hyperparameters $\boldsymbol{\eta}$. This structure naturally arises in many domains with inherent hierarchies, such as patients in pharmacokinetic models [185], longitudinal patient omics [91], single-cells in gene expression models [109], or participants in trails in cognitive modeling [100].
@@
Such models allow individual parameters to be informed by global estimates (37), stabilizing inference at low $N_{r}$ by borrowing information across observations. Even though partial pooling has been espoused as a default choice in Bayesian analysis [56, 115], it typically incurs high computational cost and has only recently received attention in SBI [70, 65, 5, 20].
To amortize two-level models, for example, we can train two diffusion models, one for the global and one for the local posterior. During training, the local model is conditioned on the true parameter $\boldsymbol{\eta}$ and during inference, an ancestral sampling scheme can be used by sampling first from the global and then from the local model. Both models can share an embedding for individual observations. However, targeting the joint posterior (35) or even the global posterior (36) without composition would require $R$ simulations per training instance. Thus, [4] ported compositional score estimation to the hierarchical setting, enabling amortized hierarchical estimation without ever simulating the full hierarchical model. To address stability issues, the authors applied an error-damping schedule $d(t)$ to the reverse diffusion with a mini-batch sampler using $M\ll S$ randomly sampled observations at each time step $t$. The resulting compositional estimator,
$\nabla_{\boldsymbol{\eta}_{t}}\!\log p(\boldsymbol{\eta}_{t}\mid\{\mathbf{y}_{\text{obs}}^{(r)}\}_{r=1}^{R})=d(t)\bigg((1-R)(1-t)\nabla_{\boldsymbol{\eta}}\!\log p(\boldsymbol{\eta})\bigl\rvert_{\boldsymbol{\eta}=\boldsymbol{\eta}_{t}}+\frac{R}{M}\sum_{m=1}^{M}\nabla_{\boldsymbol{\eta}_{t}}\!\log p(\boldsymbol{\eta}_{t}\mid\mathbf{y}_{\text{obs}}^{(m)})\bigg),$ (38)